Induction 3. In class we discussed telescoping series, meaning series of the forrm In class, we said that the partial sums
Answer #1

\sum_{n=1}^{\infty}(b_n-b_{n+1})

S_1=b_1-b_2 ----(2)

S_{n+1}=S_n+(b_{n+1}-b_{n+2}) ----(3)

n=1 (putting in equation(3))

S_{2}=S_1+(b_{2}-b_{3})

S_{2}=(b_1-b_2)+(b_{2}-b_{3}) (using value of S1 from equation(2)

S_{2}=b_1-b_{3}

n=2

S_{3}=S_2+(b_{3}-b_{4})

S_{3}=(b_1 -b_3)+(b_{3}-b_{4}) (from above S2=b1-b3)

S_{3}=b_1 -b_{4}

n=3

S_{4}=S_3+(b_{4}-b_{5})

S_{4}=(b_1 -b_4)+(b_{4}-b_{5})

S_{4}=b_1 -b_{5}

n=4

S_{5}=S_4+(b_{5}-b_{6})

S_{5}=(b_1-b_5)+(b_{5}-b_{6})

S_{5}=b_1-b_{6}

n=5

S_{6}=S_5+(b_{6}-b_{7})

S_{6}=(b_1-b_6)+(b_{6}-b_{7})

S_{6}=b_1-b_{7}

.

.

.

So we can conclude from above

S_{n}=b_1-b_{n+1}